OpenStudy (anonymous):

If you have a block that is 2.0kg being pulled horizontal to the surface and parallel to the surface. The coefficient of friction is 0.30. a) What is the acceleration? b) What is the velocity after traveling for 8 seconds? c) What is the position after traveling for 8 seconds?

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

use Newtons 2nd law and compare it with the friction force stuff.

OpenStudy (amistre64):

F = ma F = un ma = un a = un/m

OpenStudy (anonymous):

In the Force Diagram if I would to make one, would it be equal to the Force of Gravity?

OpenStudy (vincent-lyon.fr):

It depends on the force applied. There must be some missing information in your question.

OpenStudy (amistre64):

veolcity and position are then just antiderivative constructs of acceleration

OpenStudy (anonymous):

Ahhh yea sorry applied force is 10.0N

OpenStudy (vincent-lyon.fr):

Try to draw a sketch of the situation with all the forces at stake.

OpenStudy (anonymous):

This is the Force Diagram correct? 

OpenStudy (vincent-lyon.fr):

Correct, but no use to have negative values of forces on your sketch. Use Fk = µ.Fn = µ.m.g

OpenStudy (vincent-lyon.fr):

If not stated otherwise, use g = 10 m/s²

OpenStudy (anonymous):

I know I was just generalizing them: So then Fk would be 5.9N and negative due to the direction it is in?

OpenStudy (anonymous):

Then sum of Forces in x direction is 10.0N + (-5.9N) = 4.1

OpenStudy (vincent-lyon.fr):

Ok

OpenStudy (anonymous):

Then since F = ma 4.1N = (2.0kg)(a) a = 2.1m/s^2

OpenStudy (vincent-lyon.fr):

Correct

OpenStudy (anonymous):

that is part a part b would be using v = a(dt) so v = (2.1m/s^2)(8s) = 17 m/s

OpenStudy (anonymous):

right?

OpenStudy (vincent-lyon.fr):

Yep! if initial velocity is zero.

OpenStudy (anonymous):

And position is x = .5at^2 so x = .5(2.1m/s^2)(8s)^2 = 67m

OpenStudy (vincent-lyon.fr):

Yep!

OpenStudy (vincent-lyon.fr):

Well done!

OpenStudy (anonymous):

Ok thanks! Justn eeded help on first step was all

OpenStudy (vincent-lyon.fr):

Remember slipping friction force is always equal to normal force times kinetic friction coefficient.

OpenStudy (anonymous):

Yea the part i was questioning was if Normal Force was equal and opposite of Force of Gravity

OpenStudy (anonymous):

Can you help me with a few other things?

OpenStudy (vincent-lyon.fr):

It was, because the contact is horizontal and no other vertical force is acting on the body. In different words : vertical acceleration is zero, so net vertical force is zero, so weight and normal force are opposed.

OpenStudy (vincent-lyon.fr):

I can help but only a few minutes (must go and sleep).

OpenStudy (anonymous):

Can you help me doug?